Solve the following equation: log2 x = 9logx 2


Answer by scott8148(6628)

log(x) / log(2) = 9 [log(2) / log(x)]

[log(x)]^2 = 9 [log(2)]^2

log(x) = ±3 log(2)

x = 2^3 = 8

x = 2^(-3) = 1/8